OpenStudy (sneakysquash):

Please Evaluate- 9[y^3(z-6)-5] for y=2 and z=8

alones (alones):

when you convert this\[9[2^3(8-6)-5] \] Now can you solve this, or you need help with it?

OpenStudy (alivejeremy):

Can you write the question right. .-.

OpenStudy (sneakysquash):

Hold on, give me a second to work it out...

OpenStudy (sneakysquash):

Yeah man, can you helnot out?

OpenStudy (sneakysquash):

Like walk me threw?

alones (alones):

Alright you asking me or jeremy?

OpenStudy (sneakysquash):

You preferably

OpenStudy (alivejeremy):

Is this the question 9(y^3(z−6)−5)?

OpenStudy (sneakysquash):

Yeah.

OpenStudy (sneakysquash):

You have to plug in the numbers still though

OpenStudy (alivejeremy):

9(y^3(z−6)−5) =(9)(y^3(z−6)+−5) =(9)(y^3(z−6))+(9)(−5) =9y^3z−54y^3−45 =9y^3z−54y^3−45

OpenStudy (alivejeremy):

What I got.

alones (alones):

\[\ (9)((2^3)(8−6)+−5)\] \[\ (9)((2^3)(8−6))+(9)(−5)\] \[\ 144−45 \] So your answer should be \(\ 99\)

alones (alones):

@alivejeremy seems like you copied and pasted it..

OpenStudy (sneakysquash):

Thanks guys

OpenStudy (alivejeremy):

No.

OpenStudy (alivejeremy):

I was bout to finish but you did.

alones (alones):

Well you wrote it fast.

OpenStudy (alivejeremy):

No I was already writing.

OpenStudy (sneakysquash):

Truuu!

alones (alones):

What's true?

OpenStudy (sneakysquash):

That he copy cated

OpenStudy (alivejeremy):

I already solve the question before I respond.

OpenStudy (sneakysquash):

He wrote it super fast so, yeah.

alones (alones):

Anyways, @sneakysquash now you understand?

OpenStudy (sneakysquash):

Yeah I do, thanks.

OpenStudy (alivejeremy):

That's only cuz I already had solve it.

alones (alones):

Jeremy no need to make up the exuces.

OpenStudy (alivejeremy):

I'm not.

alones (alones):

*excuses

OpenStudy (alivejeremy):

Oh god.... f off...fr

OpenStudy (sneakysquash):

Hey alones I tried to fan you but Ig I'm blocked?

alones (alones):

Anyways sneaky hope that helped (: